# Brownian motion of a single microparticle partially engulfed by a giant lipid vesicle

### v1

@@ V.2 Synthesis of silica microparticles
Silica microspheres (radius $a=0.60\pm 0.05\,\mathrm{\mu m}$) were synthesized according to the protocol described by Yu et al [36]. Briefly, in a 250 mL semibatch chemical reactor, 65 mL of ethanol, 6.75 mL of water and 0.017 g of KCl were mixed at 300 RPM. A solution composed of 33.3 mL of ethanol and 2 g of TEOS was delivered in this solution at a flow rate of 0.2 mL/min using a syringe pump. The temperature was maintained at 30°C through the reaction.
@@ V.6 Optical tweezers setup
V.7 Prepatation of lipid-coated aqueous droplets in oil
@@ Acknowledgements.
We acknowledge the support from the Doctoral College of Physics and Chemistry-Physics of Strasbourg, Agence Nationale de la Recherche EDEM (ANR-21-CE06-0042), and ITI HiFunMat at the University of Strasbourg.

### v2

@@ V.2 Synthesis of silica microparticles
Silica microspheres (radius $a=0.60\pm 0.05\,\mathrm{\mu m}$) were synthesized according to the protocol described by Yu et al [44]. Briefly, in a 250 mL semibatch chemical reactor, 65 mL of ethanol, 6.75 mL of water and 0.017 g of KCl are mixed at 300 RPM. A solution composed of 33.3 mL of ethanol and 2 g of TEOS is delivered in this solution at a flow rate of 0.2 mL/min using a syringe pump. The temperature is maintained at 30°C through the reaction. For the surface functionalization of silica microparticles with amine groups, 10 mg of the powder silica particles are dispersed in toluene followed by the addition of 60 µL of APTES. The mixture is left under agitation and reflux for 3 hours. After this step the silica microparticles are collected by centrifugation (5 minutes, 5000 RCF), and washed (3 times) with ethanol and water.
@@ V.6 Optical tweezers setup
V.7 Preparation of lipid-coated aqueous droplets in oil
@@ Acknowledgements.
We acknowledge the support of the Agence Nationale de la Recherche EDEM (ANR-21-CE06-0042), the Doctoral College of Physics and Chemistry-Physics of Strasbourg, and ITI HiFunMat at the University of Strasbourg.